Use logarithms to solve the problem.
I need to find the interest rate needed for an investment of $4200 to double in 5 years if interest is compounded continuously. (Round your answer to two decimal places.)
Thank you!

Answer by lwsshak3(11628)

I need to find the interest rate needed for an investment of $4200 to double in 5 years if interest is compounded continuously.
**
Formula for continuous compounding:
A=Pe^rt, P=initial investment, r=interest rate, t=number of years, A=amt after t years
A/P=e^rt
2=e^5r
take log of both sides
ln2=5r*lne
lne=1 (log of base=1)
5r=ln2
r=ln2/5
r≈0.1386
interest rate needed for an investment of $4200 to double in 5 years if interest is compounded continuously≈13.86%